OpenStudy (anonymous):

Can Anyone Help Me.?! Arrange the list of numbers in increasing order. 3/5, 5/9,4/7

OpenStudy (darthvader2900):

I will help you if you help me with english

OpenStudy (anonymous):

Okay.

Mehek (mehek14):

you have to turn all of those fractions into decimals

Mehek (mehek14):

\(\dfrac{3}{5}=\) \(\dfrac{5}{9}=\) \(\dfrac{4}{7}=\)